A molecule having a molar mass of 151.0 g/mol crystallizes to a cubic solid where each edge of the unit cell is 714.0 pm. Each unit cell contains 2 molecules. What is the bulk density of the crystal?
1
Expert's answer
2019-09-13T05:10:08-0400

p=m/V

m=2 x 151/6.02e23=2.5083e-22

V=714^3 pm=(7.14e-8)3

m= 2.5083e-22 / (7.14e-8)^3=1.37821133 g/sm^3